OpenStudy (anonymous):

3*|3sqrt3-2sqrt7| / (3sqrt3-2sqrt7)

OpenStudy (asnaseer):

This is similar to your previous question. Basically you just need to determine if:\[3\sqrt{3}\gt2\sqrt{7}\]

OpenStudy (ash2326):

We have \[\large \frac{3*|3 \sqrt3-2 \sqrt7|}{ (3\sqrt 3-2 \sqrt7)}\] Let's check if \[3 \sqrt 3>2 \sqrt 7\] square both sides \[9 \times 7> 4 \times 7\] yeah it's greater so Can you solve it now?

OpenStudy (anonymous):

I think so. Thanks!

OpenStudy (anonymous):

@@ash2326 I don't understand your post... :( when you square them... \[\Large 3\sqrt3 > 2\sqrt7 \quad \longrightarrow \quad 9\times \underbrace{7}_{?} > 4\times 7 ??\]

OpenStudy (anonymous):

@ash2326

OpenStudy (ash2326):

Sorry I exchanged them it'd be then \[3 \sqrt 3< 2 \sqrt 7\] so \[27<28\] so @erkot I'm sorry